Giselle pays $190 in advance on her account at the athletic club.Each time she uses the club,$5 is deducted from the account.The situation can be moldeled by the equation b=190-5x,where x is the number of visits and b is the total account balance
graph the equation and find the account balance after 7 visits

Found 2 solutions by mananth, tbone streeter:
Answer by mananth(16946)

b= 190-5x
x= 7
b= 190-5*7
b=155 will be the balance
...
m.ananth@hotmail.ca

Answer by tbone streeter(1)

b= 190-5x
x= 7
b= 190-5*7
b=155 will be the balance